### v2

# On the fixed locus of the antisymplectic involution of an EPW cube

Abstract.
[…]
Key words and phrases:
Hyperkähler manifolds, antisymplectic involutions, Lagrangian subvariety, moduli spaces, Bridgeland stability, LLV decomposition.
2020 Mathematics Subject Classification:
14C20, 14D06, 14F45, 14J42, 14J50
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (ERC-2020-SyG-854361-HyperK)
1. Introduction
[…]
  • describe a family of certain singular degenerations of EPW cubes;
  • use this family and the degeneration methods of Flapan–Macrì–O’Grady–Saccà, who studied in [FMOS22] and [FMOS23] the fixed loci of antisymplectic involutions on hyper-Kähler varieties with a polarization of square 22 to compute some numerical invariants of the fixed locus.
EPW cubes are defined as follows. Let $V_{6}$ be a 6-dimensional complex vector space and consider the symplectic space ${\bigwedge\thinspace^{3}}\thinspace{V}_{6}$. For any Lagrangian subspace $A\subseteq\textstyle{\bigwedge\thinspace^{3}}\thinspace{V}_{6}$, we consider the degeneracy loci
${\mathsf{Z}}^{\geq i}_{A}\coloneqq\{[U]\in\operatorname{\mathsf{Gr}}(3,V_{6})\mid\dim(A\cap\textstyle{\bigwedge\thinspace^{2}}\thinspace{U}\wedge V_{6})\geq i\}$
in the Grassmannian $\operatorname{\mathsf{Gr}}(3,V_{6})$.
When $A$ contains no decomposable vectors, that is, when the intersection ${\bf P}(A)\cap\operatorname{\mathsf{Gr}}(3,V_{6})\subseteq{\bf P}(\textstyle{\bigwedge\thinspace^{3}}\thinspace{V}_{6})$ is empty, the stratum ${\mathsf{Z}}_{A}\coloneqq{\mathsf{Z}}^{\geq 2}_{A}$ is a normal integral 6-fold and its singular locus is the normal integral 3-fold ${\mathsf{Z}}^{\geq 3}_{A}$. Moreover, there exists a double cover
(1) $g_{A}\colon\widetilde{\mathsf{Z}}_{A}\longrightarrow{\mathsf{Z}}_{A},$
constructed in [DK20, Theorem 5.7], branched along ${\mathsf{Z}}^{\geq 3}_{A}$. The variety $\widetilde{\mathsf{Z}}_{A}$ is called an EPW cube.
[…]
In the first part of the paper (Sections 2 and 3), we study the cohomology ring $H^{\bullet}(\widetilde{\mathsf{Z}}_{A},{\bf Q})$ and the action of $\iota^{*}_{A}$ on it to prove the following theorem.
Theorem 1.1.
Let $\widetilde{\mathsf{Z}}_{A}$ be a smooth EPW cube with associated involution $\iota_{A}$. The fixed locus ${\mathsf{W}}_{A}$ of $\iota_{A}$ is a smooth irreducible threefold of general type, with canonical bundle $\omega_{{\mathsf{W}}_{A}}=\mathscr{O}_{{\mathsf{W}}_{A}}(2)$ and self-intersection number in $\widetilde{\mathsf{Z}}_{A}$ given by
${\mathsf{W}}_{A}]^{2}=\chi_{\mathrm{top}}({\mathsf{W}}_{A})=-1200.$
Our method of proof uses the action of the so-called Looijenga–Lunts–Verbitsky (LLV) algebra $\mathfrak{g}(X)$ on the cohomology ring $H^{\bullet}(X,{\bf Q})$ of a hyper-Kähler variety $X$. This action induces a decomposition, called the LLV decomposition, of $H^{\bullet}(X,{\bf Q})$ into irreducible representations of $\mathfrak{g}(X)$. One of these irreducible representations is the so-called Verbitsky component $SH(X,{\bf Q})$, defined as the subalgebra of $H^{\bullet}(X,{\bf Q})$ generated by $H^{2}(X,{\bf Q})$.
[…]
As observed in Remark 3.2, the Lagrangian submanifold ${\mathsf{W}}_{A}$ of $\widetilde{\mathsf{Z}}_{A}$ is atomic in the sense of [Bec25, Definition 1.1]. After [Mar24, Section 7.4], a natural question is to ask whether it is rigid inside $\widetilde{\mathsf{Z}}_{A}$. In the second part of the paper (Sections 4 and 5), we use the degeneration method of [FMOS22] and [FMOS23] to show that the answer is positive.
Theorem 1.2.
Let $\widetilde{\mathsf{Z}}_{A}$ be a smooth EPW cube with associated involution $\iota_{A}$. The fixed locus ${\mathsf{W}}_{A}$ of $\iota_{A}$ is a rigid atomic Lagrangian submanifold, namely its first Betti number is 0.
The degeneration we use is interesting in its own right and complements the study in [Riz24] of another family of singular EPW cubes. It is constructed as follows. Let $(S,L)$ be a polarized K3 surface of degree $2$, with associated double cover $S\to{\bf P}^{2}$ branched over a smooth sextic curve $\Gamma\subseteq{\bf P}^{2}$. The class $2L-\delta$ is big and nef on the Hilbert scheme $S^{[3]}$, and it has square $4$ and divisibility $2$. Therefore, it induces a contraction $S^{[3]}\to X$, and the involution of $S$ given by the double cover induces an involution $\iota$ on $X$.
The singular variety $X$ is a degeneration of EPW cubes. More precisely, we construct a family $\mathscr{X}\to B$, whose general fibers are (smooth) EPW cubes and whose central fiber is $X$, which carries an involution $\iota_{\mathscr{X}}$ that restricts to $\iota_{A}$ on each smooth fiber $\widetilde{\mathsf{Z}}_{A}$, and to $\iota$ on the special fiber.
Results from [BM14a] on contractions of moduli spaces of sheaves on K3 surfaces allow us to explicitly describe $X$, the involution $\iota$, and its fixed locus. The latter is a disjoint union of two components, one of dimension $2$ (which does not deform), and $F_{3}$, of dimension $3$. The normalization of the component $F_{3}$ is a birational morphism
$\Gamma^{(3)}\longrightarrow F_{3}$
which restricts to an étale double cover $E_{3}\to\operatorname{Sing}(F_{3})$ on a divisor $E_{3}$ of $\Gamma^{(3)}$.
One component of the fixed locus of $\iota_{\mathscr{X}}$ induces a family $\mathscr{Y}\to B$ of relative dimension $3$, with general fiber ${\mathsf{W}}_{A}$ and special fiber $F_{3}$. This family is not semistable, so the Clemens–Schmid method cannot be used to determine the Hodge decomposition of ${\mathsf{W}}_{A}$. However, the fibers of $\mathscr{Y}\to B$ are semi-log canonical, hence the numbers $h^{i}(\mathscr{X}_{b},\mathscr{O}_{\mathscr{X}_{b}})$ are constant. In particular, the “external” Hodge numbers of ${\mathsf{W}}_{A}$ are given by the dimension of the cohomology groups of the structure sheaf of $F_{3}$.
As in [FMOS23], the most delicate part is showing that the (schematic) central fiber of $\mathscr{Y}\to B$ is equal to $F_{3}$, and in particular proving that it is reduced. This is done in Section 4.3 where, using the Kuranishi map, we determine the local structure the fixed locus of $\iota$.
Structure of the paper
The paper is organized as follows.
[…]
In Section 3, we specialize the previous analysis to the case of EPW cubes, to obtain Theorem 1.1. We also compute the cohomology class ${\mathsf{W}}_{A}]$ (Proposition 3.1) and the Chern classes of ${\mathsf{W}}_{A}$ (Section 3.2).
In Section 4, we construct the degeneration $\mathscr{X}\to D$ and describe the special fiber $X$ and its involution.
Finally, in Section 5, we construct and study the degeneration $\mathscr{Y}\to D$, and prove that the first cohomology group $H^{1}({\mathsf{W}}_{A},{\bf Q})$ is trivial.
Acknowledgements
I would like to thank my advisor, Olivier Debarre, for suggesting me this problem and for his insights and hints. I am also grateful to Emanuele Macrì for the helpful discussions on the study of fixed loci in degenerations, to Jieao Song for discussing with me Remark 2.7 and to Mirko Mauri for pointing out Kollár’s result [Kou93, Corollary 2.64]. I thank Pietro Beri for useful discussions, and Francesco Denisi, Giovanni Mongardi, and Vanja Zuliani for their interest in this work.
2. Hyper-Kähler manifolds of $\operatorname{{K3}}^{[3]}$-type
Let $X$ be a hyper-Kähler manifold of $\operatorname{{K3}}^{[3]}$-type. Its
nonzero Hodge numbers are
${b_{0}=1}$${1}$${b_{2}=23}$${1}$${21}$${1}$${b_{4}=299}$${1}$${22}$${253}$${22}$${1}$${b_{6}=2554}$  ${1}$${21}$${253}$${2004}$${253}$${21}$${1}$
We denote by $\operatorname{SH}^{2\bullet}(X,{\bf Q})$ the image of the canonical map
$\operatorname{Sym}^{\bullet}\!H^{2}(X,{\bf Q})\longrightarrow H^{2\bullet}(X,{\bf Q})$
(the so-called Verbitsky component). Let $c_{2},c_{4},c_{6}$ be the nonzero Chern classes of $X$. One has ([Mar20, Lemma 1.5])
$c_{2}\in\operatorname{Sym}^{2}\!H^{2}(X,{\bf Q})=\operatorname{SH}^{4}(X,{\bf Q})\subseteq H^{4}(X,{\bf Q}).$
The nonzero Chern numbers are ([EGL01, Theorem 0.1])
(2) $c_{6}=3200\ ,\ c_{2}c_{4}=14720\ ,\ c_{2}^{3}=36800.$
Recall that, given a class $\alpha\in H^{4k}(X,{\bf Q})$ that remains of type $(2k,2k)$ on all small deformations of $X$, the generalized Fujiki constant is a nonzero rational number $C(\alpha)$ that satisfies
(3) $\forall\beta\in H^{2}(X,{\bf Q})\qquad\int_{X}\alpha\cdot\beta^{6-2k}=C(\alpha)q_{X}(\beta)^{3-k}.$
From [Son22, Lemma 2.1.7], we have [^1] All Chern numbers and Fujiki constants for hyper-Kähler manifolds of $\operatorname{{K3}}^{[n]}$-type can be computed using https://github.com/8d1h/bott.
(4) $C(1)=15,\quad C(c_{2})=108,\quad C(c_{2}^{2})=1200,\quad C(c_{4})=480.$
Polarizing both terms of Equation (3), we obtain that for all $\beta_{1},\dots,\beta_{6-2k}\in H^{2}(X,{\bf Q})$,
(5) $\int_{X}\alpha\beta_{1}\cdots\beta_{6-2k}=\frac{C(\alpha)}{(6-2k)!}\sum_{\sigma\in\mathfrak{S}_{6-2k}}q(\beta_{\sigma(1)},\beta_{\sigma(2)})\cdots q(\beta_{\sigma(6-2k-1)},\beta_{\sigma(6-2k)}),$
where $\mathfrak{S}_{6-2k}$ is the group of permutations of 62k6-2k elements.
[…]
Following [GKLR22, Section 2], we denote by $\bar{V}$ the vector space $H^{2}(X,{\bf Q})$ with the (rational) Beauville–Bogomolov–Fujiki form $q$, and by $V$ the Mukai completion
(6) $V=H^{2}(X,{\bf Q})\oplus U$
with the quadratic form $q\oplus\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$, so that $U\simeq{\bf Q}^{\oplus 2}$ with the structure of a hyperbolic plane.
[…]
(7) $H^{\bullet}(X,{\bf Q})=V_{(3)}\oplus V_{(1,1)},$
where $V_{(3)}$ is the Verbitsky component $\operatorname{SH}^{2\bullet}(X,{\bf Q})$ and $V_{(1,1)}$ is ${\bigwedge\thinspace^{2}}\thinspace{V}$. Furthermore, under the action of $\mathfrak{so}(\bar{V})$, the cohomology groups decompose further into the following representations
$\begin{array}[]{rcl}H^{0}(X,{\bf Q})&=&{\bf Q},\\ H^{2}(X,{\bf Q})&=&\bar{V},\\ H^{4}(X,{\bf Q})&=&\operatorname{Sym}^{2}\!\bar{V}\oplus\bar{V},\\ H^{6}(X,{\bf Q})&=&\operatorname{Sym}^{3}\!\bar{V}\oplus\textstyle{\bigwedge\thinspace^{2}}\thinspace{\bar{V}}\oplus{\bf Q}\eta,\\ H^{8}(X,{\bf Q})&=&\operatorname{Sym}^{2}\!\bar{V}\oplus\bar{V},\\ H^{10}(X,{\bf Q})&=&\bar{V},\\ H^{12}(X,{\bf Q})&=&{\bf Q},\end{array}$
where the summands in the fourth line are pairwise orthogonal with respect to the intersection pairing. [^3] The class $\eta$ was studied in [HHT12, Proposition 6.1]. There are further decompositions
$\operatorname{Sym}^{2}\!\bar{V}$ $=\bar{V}_{(2)}\oplus{\bf Q},$
$\operatorname{Sym}^{3}\!\bar{V}$ $=\bar{V}_{(3)}\oplus V,$
into irreducible representations, whereas ${\bigwedge\thinspace^{2}}\thinspace{\bar{V}}=\bar{V}_{(1,1)}$ is already irreducible.
For $X$ very general (so nonprojective), the (special) Mumford–Tate algebra of $X$ is equal to the algebra $\mathfrak{so}(\bar{V})$ ([GKLR22, Proposition 2.38]). Therefore the space of Hodge classes in $H^{\bullet}(X,{\bf Q})$ coincides with its $\mathfrak{so}(\bar{V})$-invariant part ([GKLR22, Section 2.3]): it is given by
(8) $\begin{array}[]{rcl}\operatorname{Hdg}^{0}(X,{\bf Q})&=&{\bf Q},\\ \operatorname{Hdg}^{2}(X,{\bf Q})&=&0,\\ \operatorname{Hdg}^{4}(X,{\bf Q})&=&{\bf Q}c_{2},\\ \operatorname{Hdg}^{6}(X,{\bf Q})&=&{\bf Q}\eta,\\ \operatorname{Hdg}^{8}(X,{\bf Q})&=&{\bf Q}c_{4},\\ \operatorname{Hdg}^{10}(X,{\bf Q})&=&0,\\ \operatorname{Hdg}^{12}(X,{\bf Q})&=&{\bf Q}.\end{array}$
In particular, the classes $c_{4}$ and $c_{2}^{2}$ are proportional and the proportionality coefficient can be determined using the constants given in (3): one finds
(9) $c_{2}^{2}=\frac{C(c_{2}^{2})}{C(c_{4})}\,c_{4}=\frac{1200}{480}\,c_{4}=\frac{5}{2}c_{4}.$
[…]
2.2. Action of automorphisms on the LLV algebra
Given an automorphism $\varphi$ of finite order of $X$, we want to study the induced automorphism $\varphi^{*}$ of $H^{\bullet}(X,{\bf Q})$.
The automorphism $\varphi^{*}_{2}$ of $\bar{V}=H^{2}(X,{\bf Q})$ extends to an automorphism $\gamma$ of $V$ by acting as the identity on the hyperbolic plane $U$ (see Equation (6)). This induces automorphisms $\gamma_{\lambda}$ of any $\mathfrak{g}$-representation $V_{\lambda}$. We compare in Lemma 2.2 the automorphism $\varphi^{*}$ with the automorphism $\gamma_{(3)}\oplus\gamma_{(1,1)}$ of $H^{\bullet}(X,{\bf Q})$ obtained from the isomorphism (7). In Lemma 3.1, we show that the two actions are equal for the antisymplectic involution associated to EPW cubes.
Lemma 2.1.
Let $X$ be a hyper-Kähler manifold and let $\varphi$ be an automorphism of $X$. For each $\mathfrak{g}$-stable factor $W$ of $H^{\bullet}(X,{\bf Q})$, $\varphi^{*}(W)$ is again $\mathfrak{g}$-stable and isomorphic to $W$.
Proof.
By [Flo22, Lemma 2.3], the application $h\mapsto(\varphi^{*})h(\varphi^{*})^{-1}$ induces an automorphism of $\mathfrak{g}$. In particular, $\mathfrak{g}\varphi^{*}(W)=\varphi^{*}\mathfrak{g}(W)=\varphi^{*}(W)$ for any $\mathfrak{g}$-invariant subspace $W$. ∎
The lemma implies that, in our case, the automorphism $\varphi^{*}$ of $H^{\bullet}(X,{\bf Q})$ preserves the irreducible components $V_{(3)}$ and $V_{(1,1)}$, hence it decomposes as $\varphi^{*}=\varphi^{*}_{(3)}\oplus\varphi^{*}_{(1,1)}$.
Since the map from the Verbitsky component to $H^{*}(X,{\bf Q})$ is given by cup product, one has $\varphi^{*}_{(3)}=\gamma_{(3)}$. For $\varphi^{*}_{(1,1)}$, the same argument as in [BGGG25, Proposition 3.3] shows the following.
Lemma 2.2.
Let $\varphi\in\operatorname{Aut}(X)$ be an automorphism of order $n$ and let $\varphi^{*}_{(1,1)}$ be the induced automorphism of $V_{(1,1)}$. If $n$ is odd, $\varphi^{*}_{(1,1)}=\gamma_{(1,1)}$, while if $n$ is even, $\varphi^{*}_{(1,1)}=\pm\gamma_{(1,1)}$.
Proof.
We follow the proof of [BGGG25, Proposition 3.3], borrowing its notation. Since $\varphi^{*}_{(1,1)}$ with $\gamma_{(1,1)}$ are both automorphisms of $V_{(1,1)}$, they satisfy
$\varphi^{*}_{(1,1)}(L_{\alpha}\cdot x)=L_{\varphi^{*}(\alpha)}\cdot\varphi^{*}_{(1,1)}(x),\quad\text{and}\quad\gamma_{(1,1)}(L_{\alpha}\cdot x)=L_{\varphi^{*}(\alpha)}\cdot\gamma_{(1,1)}(x)$
for all $\alpha\in H^{2}(X,{\bf Z})$ that define a $\mathfrak{sl}_{2}$-triple. In particular,
$\varphi^{*}_{(1,1)}\circ\gamma_{(1,1)}^{n-1}(L_{\alpha}\cdot x)=\varphi^{*}_{(1,1)}(L_{(\varphi^{*})^{n-1}(\alpha)}\cdot\gamma_{(1,1)}^{n-1}(x))=L_{\alpha}\cdot\varphi^{*}_{(1,1)}\circ\gamma_{(1,1)}^{n-1}(x).$
Therefore $\varphi^{*}_{(1,1)}\circ\gamma_{(1,1)}^{n-1}$ is a $\mathfrak{g}$-isomorphism of $V_{(1,1)}$, hence, it is equal to $\lambda\cdot\operatorname{Id}$ by Schur’s lemma, for some $\lambda\in{\bf Q}$. Comparing orders, we obtain $\lambda^{n}=\operatorname{Id}$, hence $\lambda$ is $1$ for $n$ odd, and equal to $\pm 1$ for even $n$. ∎
[…]
In this case, the (special) Mumford–Tate algebra of $X$ is equal to $\mathfrak{so}(\bar{T})$, and the $\mathfrak{so}(\bar{T})$-invariant part of $H^{\bullet}(X,{\bf Q})$ is the space of Hodge classes.
One has $\operatorname{Hdg}^{2}(X,{\bf Q})={\bf Q}h$.
The space $\operatorname{Hdg}^{4}(X,{\bf Q})$ has dimension 3 (recall that the invariant part of $\operatorname{Sym}^{2}\bar{T}$ has dimension 1). We denote by $\lambda$ a generator of the trivial factor outside the Verbitsky component.
The space $\operatorname{Hdg}^{6}(X,{\bf Q})$ also has dimension 3 and intersects the Verbitsky component in dimension 22. Up to multiplication by a nonzero rational number, we obtain that $h\lambda$ is the absolute Hodge class $\eta$ from (8). By [HHT12, Proposition 6.1], its square is $4$ (modulo squares of nonzero rational numbers).
The following lemma gives us a basis of $\mathrm{Hdg}^{\bullet}(X,{\bf Q})$.
Lemma 2.3.
The families $(h^{2},c_{2})$ in $\operatorname{Hdg}^{4}(X,{\bf Q})$ and $(h^{3},hc_{2})$ in $\operatorname{Hdg}^{6}(X,{\bf Q})$ are linearly independent.
Proof.
The statement in degree $4$ follows from the one in degree $6$, since any nontrivial linear relation between $h^{2}$ and $c_{2}$ would give a nontrivial linear relation between $h^{3}$ and $hc_{2}$.
We show that $h^{3}$ and $hc_{2}$ are linearly independent. From Equation (5), we get that there exist nonzero constants $x$, $y$, and $z$ such that
$\int_{X}h^{3}\beta^{3}$ $=xq(h)q(\beta)q(h,\beta)+yq(h,\beta)^{3},$
$\int_{X}hc_{2}\beta^{3}$ $=zq(h,\beta)q(\beta)$
for all $\beta\in H^{2}(X,{\bf Z})$. Assume that the classes $h^{3}$ and $hc_{2}$ are proportional. Then for any $\beta$ such that $q(\beta)=0$, we have $q(h,\beta)=0$, which is absurd since $q$ is nondegenerate, hence the quadric $q(\beta)=0$ is not contained in any hyperplane. ∎
[…]
$\operatorname{Hdg}^{0}(X,{\bf Q})$ $={\bf Q},$
$\operatorname{Hdg}^{2}(X,{\bf Q})$ $={\bf Q}h,$
$\operatorname{Hdg}^{4}(X,{\bf Q})$ $={\bf Q}h^{2}\oplus{\bf Q}c_{2}\oplus{\bf Q}\lambda,$
(10) $\operatorname{Hdg}^{6}(X,{\bf Q})$ $={\bf Q}h^{3}\oplus{\bf Q}hc_{2}\oplus{\bf Q}h\lambda,$
$\operatorname{Hdg}^{8}(X,{\bf Q})$ $={\bf Q}h^{4}\oplus{\bf Q}h^{2}c_{2}\oplus{\bf Q}h^{2}\lambda,$
$\operatorname{Hdg}^{10}(X,{\bf Q})$ $={\bf Q}h^{5},$
$\operatorname{Hdg}^{12}(X,{\bf Q})$ $={\bf Q}h^{6},$
where the class $\lambda$ depends on the polarization $h$, while the class $h\lambda$ is the absolute Hodge class $\eta$, and is orthogonal to the Verbitsky component $\operatorname{SH}^{3}(X,{\bf Q})$.
The intersection $\operatorname{Hdg}^{\bullet}(X,{\bf Q})\cap\operatorname{SH}^{2\bullet}(X,{\bf Q})$ is the $\bf Q$}-subalgebra of $H^{\bullet}(X,{\bf Q})$ generated by $\{h,c_{2},c_{4},c_{6}\}$. We determine the relations in each degree.
Proposition 2.4.
(a) There are no relations in degree $\leq 6$.
(b) In degree $8$, the relations are
$c_{4}=-\frac{160}{q(h)^{2}}h^{4}+\frac{80}{3q(h)}h^{2}c_{2},\quad c_{2}^{2}=\frac{5}{2}c_{4}.$
(c) In degree $10$, the relations are
$h^{3}c_{2}=\frac{36}{5q(h)}\,h^{5},\quad hc_{2}^{2}=\frac{80}{q(h)^{2}}\,h^{5},\quad hc_{4}=\frac{32}{q(h)^{2}}\,h^{5}.$
(d) In degree $12$, we have
$h^{6}=15\,q(h)^{3},\quad h^{4}c_{2}=108\,q(h)^{2},\quad h^{2}c_{2}^{2}=1200\,q(h),\quad h^{2}c_{4}=480\,q(h).$
Proof.
The relations in degree $12$ follow from the values of the generalized Fujiki constants given in (4) and their definition (3). The relations in degree $10$ then follow from the fact that the vector space $\operatorname{Hdg}^{10}(X,{\bf Q})$ has dimension one. For example, we obtain
$h^{3}c_{2}=\frac{h^{4}c_{2}}{h^{6}}\,h^{5}=\frac{108\,q(h)^{2}}{15\,q(h)^{3}}\,h^{5}=\frac{36}{5q(h)}\,h^{5}$
and similarly for the other relations.
In degree $8$, the second relation is (9). For the first relation, write $c_{4}=xh^{4}+yh^{2}c_{2}$ for some $x,y\in{\bf Q}$. Taking the product with $h^{2}$, we get
$480\,q(h)=15\,xq(h)^{3}+108\,yq(h)^{2}.$
Taking the product with $c_{2}$ and using the Chern numbers given in (2), we get
$14720=108\,xq(h)^{2}+1200\,yq(h).$
Solving this system of equations for $x$ and $y$, we obtain the desired relation. ∎
Remark 2.5.
One also have the following relations involving $\lambda$. The class $\eta=h\lambda$ is orthogonal to the Verbitsky component $\operatorname{SH}^{3}(X,{\bf Q})$ and (up to multiply $\eta$ by a rational number), we can suppose $\eta^{2}=4$. Namely, in degree $12$ we have the following relations
$h^{2}\lambda^{2}=4,\qquad h^{4}\lambda=0,\qquad h^{2}c_{2}\lambda=0,\qquad c_{2}^{2}\lambda=0,$
where the last equality is obtain from the fact that $c_{2}^{2}$ is a linear combination of $h^{4}$ and $h^{2}c_{2}$.
Since $\operatorname{Hdg}^{5}(X)={\bf Q}\cdot h^{5}$, the above equalities together with Proposition 2.4 imply
$h\lambda^{2}=\frac{4}{15q(h)^{3}}h^{5},\qquad h^{3}\lambda=0,\qquad hc_{2}\lambda=0.$
We conclude this section by studying the class of Lagrangian submanifolds of hyper-Kähler manifolds of $\operatorname{{K3}}^{[3]}$}-type.
Proposition 2.6.
Let $W$ be a Lagrangian submanifold of a very general polarized hyper-Kähler manifold $(X,h)$ of $\operatorname{{K3}}^{[3]}$}-type. The orthogonal projection to the Verbitsky component $\mathrm{SH}^{6}(X,{\bf Q})$ of the class $W]\in H^{6}(X,{\bf Q})$ is equal to
$\overline{[W]}\coloneqq\frac{[W]\cdot h^{3}}{72\cdot q(h)^{2}}\left(\frac{12}{q(h)}h^{3}-hc_{2}\right).$
Proof.
The class $[W]$ is a 3-dimensional Hodge class hence, by Equation (2.3), it can be written as
$W]=ah^{3}+bhc_{2}+ch\lambda,$
where $a,b,c\in{\bf Q}$, and $h\lambda$ is orthogonal to the Verbitsky component. In particular, $\overline{[W]}=ah^{3}+bhc_{2}$ and, for any class $\gamma\in\operatorname{SH}^{6}(X,{\bf Q})$, we have $\gamma\cdot[W]=\gamma\cdot\overline{[W]}$.
We apply this to the class $\gamma=h\sigma\bar{\sigma}$ in $\operatorname{SH}^{6}(X,{\bf Q})$, where $\sigma$ is a symplectic form on $X$. Since $W$ is Lagrangian, it satisfies $\gamma\cdot[W]=0$ and thus $\gamma\cdot\overline{[W]}=0$. Using Equation (5), the Fujiki constants computed in Equation (4), and the fact that $q(\sigma,h)=0$, we compute
$\gamma\cdot\overline{[W]}$ $=a\int_{X}h^{4}\sigma\bar{\sigma}+b\int_{X}c_{2}h^{2}\sigma\bar{\sigma}$
$=a\frac{C(1)}{6!}6\cdot 4!q(h)^{2}q(\sigma,\bar{\sigma})+b\frac{C(c_{2})}{4!}8q(h)q(\sigma,\bar{\sigma})$
$=q(h)q(\sigma,\bar{\sigma})(3q(h)a+36b).$
Since $\gamma\cdot\overline{[W]}=0$ and $q(h)q(\sigma,\bar{\sigma})>0$, we obtain $a=-\frac{12}{q(h)}b$.
Finally, we determine $b$ in terms of the degree of $W$. Indeed, we have
$W]\cdot h^{3}=\overline{[W]}\cdot h^{3}=b\Bigl(hc_{2}-\frac{12}{q(h)}h^{3}\Bigr)\cdot h^{3}=-72\,q(h)^{2}b,$
where, in the last equality, we use the numbers computed in Proposition 2.4. This gives the value of $b$ and proves the proposition. ∎
Remark 2.7.
Jieao Song pointed out to me that the above formula generalizes to any polarized hyper-Kähler variety $X$ of dimension $2n$ with polarization $h$, and any Lagrangian subvariety $F$ deforming with the polarization. In particular, following the same strategy of [Son22, Theorem 2.2.4], he shows that the orthogonal projection to the Verbitsky component $\operatorname{SH}^{2n}(X,{\bf Q})$ of the class $F]\in H^{2n}(X,{\bf Q})$ is equal to
(11) $\overline{[F]}=\frac{[F]\cdot h^{n}}{q(h)^{n}}\cdot\frac{\mu^{n}}{c_{X}}\overline{[\exp(h/\mu)\operatorname{td}_{X}^{1/2}]_{n}},\quad\text{where $\mu\coloneqq\sqrt{\frac{-q(h)}{2r_{X}}}$.}$
Here $c_{X}=\frac{C(1)}{(2n-1)!!}$ and $r_{X}\coloneqq\frac{C(c_{2})2^{n}n!(2n-1)}{(2n)!24C(1)}$ is the constant introduced in [Bec23, Equation (3.1)].
The key point is to generalize [Son22, Lemma 2.2.1] by showing that for any class $\alpha\in H^{2}(X,{\bf Q})$ there is an equality
$(\iota^{*}\alpha)^{n}\cdot[F]=\frac{[F]\cdot h^{n}}{q(h)^{n}}q(\alpha,L)^{n}.$
By writing $\operatorname{td}_{X}^{1/2}=\left(\operatorname{td}^{1/2}_{2k}\right)_{k}$ where $\operatorname{td}_{2k}^{1/2}\in H^{4k}(X,{\bf Q})$ and developing Equation (11), one obtain
(12) $\overline{[F]}=\frac{[F]\cdot h^{n}}{q(h)^{n}\cdot c_{X}}\sum_{2k\leq n}\left(\frac{-q(h)}{2r_{X}}\right)^{k}\frac{h^{n-2k}}{(n-2k)!}\ \overline{\operatorname{td}_{2k}^{1/2}}.$
[…]
$\overline{[F]}=\frac{[F]\cdot h^{3}}{q(h)^{3}}\left(\frac{h^{3}}{3!}-\frac{q(h)}{3\cdot 24}hc_{2}\right),$
which is exactly Proposition 2.6.
3. EPW cubes
Let $A\subseteq\textstyle{\bigwedge\thinspace^{3}}\thinspace{V}_{6}$ be a Lagrangian subspace with no decomposable vectors and such that the degeneracy locus ${\mathsf{Z}}_{A}^{\geq 4}$ is empty. The double cover $g_{A}\colon\widetilde{\mathsf{Z}}_{A}\to{\mathsf{Z}}_{A}$ defined in Equation (1) is then branched over the smooth threefold ${\mathsf{Z}}_{A}^{\geq 3}\subseteq{\mathsf{Z}}_{A}$ and the EPW cube $\widetilde{\mathsf{Z}}_{A}$ is a smooth hyper-Kähler manifold of K3[3]-type. The ample class $h\coloneqq g_{A}^{*}\mathscr{O}_{\operatorname{\mathsf{Gr}}(3,V_{6})}(1)$ satisfies $q(h)=4$.
The smooth threefold ${\mathsf{W}}_{A}\coloneqq g_{A}^{-1}({\mathsf{Z}}_{A}^{\geq 3})\subseteq\widetilde{\mathsf{Z}}_{A}$ is the fixed locus of the involution $\iota_{A}$ associated to $g_{A}$. In this section, using results and numerical relations proved in Section 2, we compute the Euler characteristic of ${\mathsf{W}}_{A}$ as well as the cohomology class ${\mathsf{W}}_{A}]\in H^{6}({\mathsf{Z}}_{A},{\bf Q})$. Additionally, we determine the Chern classes of ${\mathsf{W}}_{A}$.
3.1. Euler characteristic of ${\mathsf{W}}_{A}$
The double cover $g_{A}$ gives the equality
$\chi_{\textnormal{top}}(\widetilde{\mathsf{Z}}_{A})-\chi_{\textnormal{top}}({\mathsf{W}}_{A})=2\bigl(\chi_{\textnormal{top}}({\mathsf{Z}}_{A})-\chi_{\textnormal{top}}({\mathsf{Z}}^{\geq 3}_{A})\bigr),$
between topological Euler characteristics. Since ${\mathsf{W}}_{A}\simeq{\mathsf{Z}}^{\geq 3}_{A}$, we obtain
(13) $\chi_{\textnormal{top}}({\mathsf{W}}_{A})=2\chi_{\textnormal{top}}({\mathsf{Z}}_{A})-\chi_{\textnormal{top}}(\widetilde{\mathsf{Z}}_{A}).$
For a compact complex manifold, the topological Euler characteristic $\chi_{\textnormal{top}}$ is the top Chern class so, in particular, $\chi_{\textnormal{top}}(\widetilde{\mathsf{Z}}_{A})=3200$ by Equation (2).
The covering involution $\iota_{A}$ of $\widetilde{\mathsf{Z}}_{A}$ acts on $H^{\bullet}(\widetilde{\mathsf{Z}}_{A},{\bf Q})$ and
$H^{\bullet}({\mathsf{Z}}_{A},{\bf Q})=H^{\bullet}(\widetilde{\mathsf{Z}}_{A},{\bf Q})^{+},$
where ++ means the $\iota_{A}$-invariant part. We have
$H^{2}(\widetilde{\mathsf{Z}}_{A},{\bf Q})^{+}={\bf Q}h\quad,\quad H^{2}(\widetilde{\mathsf{Z}}_{A},{\bf Q})^{-}=h^{\perp}.$
This is because $h$ is $\iota_{A}$-invariant and the symplectic form of $\widetilde{\mathsf{Z}}_{A}$ (whose class describes, by deformation of $(\widetilde{\mathsf{Z}}_{A},h)$, a dense open subset of the intersection of $h^{\perp}$ and a smooth quadric) is $\iota_{A}$-antiinvariant.
The discussion in Section 2.2 implies that either the action of $\iota_{A}$ on $H^{\bullet}(X,{\bf Q})=V_{(3)}\oplus V_{(1,1)}$ is the action induced by the automorphism $\iota_{A}^{*}\oplus\operatorname{Id}$ of $V=H^{2}(X,{\bf Q})\oplus U$ (as above, $U$ is the hyperbolic plane); or these actions coincide on the Verbitsky component $V_{(3)}$ and are opposite on $V_{(1,1)}$.
These two possibilities for the action of $\iota_{A}$ in cohomology give two possibilities for the cohomology groups of ${\mathsf{Z}}_{A}$:
  • In the first case (natural action on $V_{(1,1)}$), we can compute
    $b_{0}({\mathsf{Z}}_{A})=b_{2}({\mathsf{Z}}_{A})=1,\quad b_{4}({\mathsf{Z}}_{A})=255,\quad b_{6}({\mathsf{Z}}_{A})=486,$
    and
    $\chi_{\textnormal{top}}({\mathsf{Z}}_{A})=2(1+1+255)+486=1000.$
  • In the second case (opposite action on $V_{(1,1)}$), we can compute
    $b_{0}({\mathsf{Z}}_{A})=b_{2}({\mathsf{Z}}_{A})=1,\quad b_{4}({\mathsf{Z}}_{A})=276,\quad b_{6}({\mathsf{Z}}_{A})=276,$
    and
    $\chi_{\textnormal{top}}({\mathsf{Z}}_{A})=2(1+1+276)+276=832.$
For example, we show how to determine $h^{6}({\mathsf{Z}}_{A},{\bf Q})$. We rewrite the decompositions in Section 2.3 as
$H^{6}(X,{\bf Q})\cap V_{(3)}=\operatorname{Sym}^{3}\!\bar{T}\oplus(\operatorname{Sym}^{2}\!\bar{T}\cdot h)\oplus(\bar{T}\cdot h^{2})\oplus{\bf Q}h^{3}$
and
$H^{6}(X,{\bf Q})\cap V_{(1,1)}=\textstyle{\bigwedge\thinspace^{2}}\thinspace{\bar{T}}\oplus(h\wedge\bar{T})\oplus\textstyle{\bigwedge\thinspace^{2}}\thinspace{U},$
where $\bar{T}\coloneqq h^{\perp}$.
The fixed subspace for the action of $\iota_{A}^{*}$ on $H^{6}(X,{\bf Q})\cap V_{(3)}$ is equal to $(\operatorname{Sym}^{2}\!\bar{T}\cdot h)\oplus{\bf Q}h^{3}$, since $\iota^{*}_{A}$ acts on $\bar{T}$ as $\operatorname{Id}$ and fixes $h$. Since $\bar{T}$ has dimension $22$, this fixed subspace has dimension 254.
We consider now the action of $\iota^{*}_{A}$ on $H^{6}(X,{\bf Q})\cap V_{(1,1)}$. By Lemma 2.2, this action is equal to $\pm\gamma_{(1,1)}$, where $\gamma_{(1,1)}$ is the action induced by the involution $\iota_{A}^{*}\oplus\operatorname{Id}$ on $V$. The fixed subspace for $\gamma_{(1,1)}$ is ${\bigwedge\thinspace^{2}}\thinspace{\bar{T}}\oplus\textstyle{\bigwedge\thinspace^{2}}\thinspace{U}$, of dimension 232, and the fixed space for $\gamma_{(1,1)}$ is equal to $h\wedge\bar{T}$, of dimension 22.
Hence, $b_{6}({\mathsf{Z}}_{A})$ is equal to $254+232=486$ in the natural case and to $254+22=276$ in the opposite case. Therefore, the relation (13) gives
(14) $\chi_{\textnormal{top}}({\mathsf{W}}_{A})=\begin{cases}-1200&\quad\text{(natural case),}\\ -1536&\quad\text{(opposite case)}.\end{cases}$
3.1.1. The class ${\mathsf{W}}_{A}]$ in $H^{6}(\widetilde{Z}_{A},{\bf Z})$
The class ${\mathsf{W}}_{A}]$ is a 3-dimensional Hodge class of $\widetilde{\mathsf{Z}}_{A}$ hence, by Equation (2.3), it can be written as
${\mathsf{W}}_{A}]=ah^{3}+bhc_{2}+c\eta,$
where $a,b,c\in{\bf Q}$, and $\eta$ is a class of square 4 (see Section 2.3).
Lemma 3.1.
The subvariety ${\mathsf{W}}_{A}\subseteq\widetilde{\mathsf{Z}}_{A}$ has class
${\mathsf{W}}_{A}]=\frac{5}{8}\left(3h^{3}-hc_{2}\right)\in H^{6}(\widetilde{\mathsf{Z}}_{A},{\bf Z}).$
Moreover, the Euler characteristic of $W_{A}$ is $\chi_{\textnormal{top}}({\mathsf{W}}_{A})=-1200$ and the action of $\iota_{A}^{*}$ in cohomology is natural.
Proof.
Since ${\mathsf{Z}}^{\geq 2}_{A}\cdot h^{3}=720$ ([IKKR19, Section 2.3]), it follows from Proposition 2.6 that the class of ${\mathsf{W}}_{A}$ in $H^{6}(\widetilde{\mathsf{Z}}^{\geq 2}_{A},{\bf Q})$ is ${\mathsf{W}}_{A}]=\frac{15}{8}h^{3}-\frac{5}{8}hc_{2}+c\eta.$ Since ${\mathsf{W}}_{A}$ is Lagrangian, it satisfies
$\chi_{\textnormal{top}}({\mathsf{W}}_{A})=[{\mathsf{W}}_{A}]^{2}=1200+4c^{2}.$
We compare this with the two possible values of $\chi_{\textnormal{top}}({\mathsf{W}}_{A})$ computed in Equation (14) to obtain that $4c^{2}$ is either 0 or $336$, hence $c=0$ and $\chi_{\textnormal{top}}({\mathsf{W}}_{A})=-1200$. ∎
Remark 3.2.
The restriction map
$H^{2}(\widetilde{\mathsf{Z}}_{A},{\bf Q})\longrightarrow H^{2}({\mathsf{W}}_{A},{\bf Q})$
has rank 1 and kernel $h^{\perp}$ (it maps the symplectic form to 0 and since the class of the latter describes a dense open subset of the intersection of $h^{\perp}$ and a smooth quadric, it vanishes on $h^{\perp}$; it is nonzero since the restriction of an ample class (such as $h$) is ample). This implies that ${\mathsf{W}}_{A}]$ is an atomic Lagrangian [Bec25, Theorem 1.8]. Notice that its class belongs to the Verbitsky component, according to what is expected in [Bec25, Proposition 3.5].
3.2. Chern classes
By the discussion after [DK20, Theorem 5.6], and [DK20, Lemma 4.3] (applied with $S=\operatorname{\mathsf{Gr}}(3,V_{6})$, $k=3$, $c_{1}(\mathscr{L})=c_{1}(\mathscr{A}_{1})=0$, and $c_{1}(\mathscr{A}_{2})=\mathscr{O}_{\operatorname{\mathsf{Gr}}(3,V_{6})}(-4)$, and giving an equality of Cartier divisor classes), we have $\omega_{{\mathsf{Z}}_{A}^{\geq 3}}=\mathscr{O}_{{\mathsf{Z}}_{A}^{\geq 3}}(2)$, so that ${\mathsf{Z}}_{A}^{\geq 3}$ (hence also ${\mathsf{W}}_{A}$) is a threefold of general type.
One can use the exact sequence
$0\to T_{{\mathsf{W}}_{A}}\to T_{\widetilde{\mathsf{Z}}_{A}}|_{{\mathsf{W}}_{A}}\to N_{{\mathsf{W}}_{A}/\widetilde{\mathsf{Z}}_{A}}\to 0$
and the isomorphism $T_{{\mathsf{W}}_{A}}\simeq N^{\vee}_{{\mathsf{W}}_{A}/\widetilde{\mathsf{Z}}_{A}}$ coming from the fact that ${\mathsf{W}}_{A}\subseteq\widetilde{\mathsf{Z}}_{A}$ is Lagrangian to deduce the relation
$c(T_{{\mathsf{W}}_{A}})c(T^{\vee}_{{\mathsf{W}}_{A}})=c(T_{\widetilde{\mathsf{Z}}_{A}})|_{{\mathsf{W}}_{A}}$
between Chern classes. Since the odd Chern classes of $\widetilde{\mathsf{Z}}_{A}$ vanish, this is equivalent to
$(-c_{1}^{2}+2c_{2})(T_{{\mathsf{W}}_{A}})=c_{2}(\widetilde{\mathsf{Z}}_{A})|_{{\mathsf{W}}_{A}}.$
Thus the Chern classes of $T_{{\mathsf{W}}_{A}}$ are
(15) $c_{1}(T_{{\mathsf{W}}_{A}})=-2h|_{{\mathsf{W}}_{A}},\quad c_{2}(T_{{\mathsf{W}}_{A}})=\frac{1}{2}c_{2}|_{{\mathsf{W}}_{A}}+2h^{2}|_{{\mathsf{W}}_{A}},\quad c_{3}(T_{{\mathsf{W}}_{A}})=\chi_{\textnormal{top}}({\mathsf{W}}_{A})=-1200.$
We can then obtain the Euler characteristics of $\mathscr{O}_{{\mathsf{W}}_{A}}$ and $\Omega_{{\mathsf{W}}_{A}}^{1}$.
Proposition 3.3.
One has $\chi({\mathsf{W}}_{A},\mathscr{O}_{{\mathsf{W}}_{A}})=-130$ and $\chi({\mathsf{W}}_{A},\Omega_{{\mathsf{W}}_{A}}^{1})=470$.
Proof.
The Hirzebruch–Riemann–Roch formula gives
$\chi({\mathsf{W}}_{A},\mathscr{O}_{{\mathsf{W}}_{A}})=\frac{1}{24}\,c_{1}(T_{{\mathsf{W}}_{A}})c_{2}(T_{{\mathsf{W}}_{A}})=-\frac{1}{24}\,(4h^{3}+hc_{2})\cdot[{\mathsf{W}}_{A}]=-130.$
Since ${\mathsf{W}}_{A}$ is a smooth projective threefold, by Poincaré duality and Hodge theory, we can write
$\frac{1}{2}\chi_{\textnormal{top}}({\mathsf{W}}_{A})$ $=1-2h^{0,1}({\mathsf{W}}_{A})+(2h^{0,2}({\mathsf{W}}_{A})+h^{1,1}({\mathsf{W}}_{A}))-(h^{0,3}({\mathsf{W}}_{A})+h^{1,2}({\mathsf{W}}_{A}))$
$=\chi({\mathsf{W}}_{A},\mathscr{O}_{{\mathsf{W}}_{A}})-h^{1,0}({\mathsf{W}}_{A})+h^{1,3}({\mathsf{W}}_{A})+h^{1,1}({\mathsf{W}}_{A})-h^{1,2}({\mathsf{W}}_{A})$
$=\chi({\mathsf{W}}_{A},\mathscr{O}_{{\mathsf{W}}_{A}})-\chi({\mathsf{W}}_{A},\Omega_{{\mathsf{W}}_{A}}^{1}),$
from which we obtain, using Lemma 3.1, the value $\chi({\mathsf{W}}_{A},\Omega_{{\mathsf{W}}_{A}}^{1})=-130-\frac{1}{2}\,(-1200)=470$. ∎
4. A singular degeneration of $\widetilde{\mathsf{Z}}^{\geq 2}_{A}$
In this section, we construct singular degenerations of $\widetilde{\mathsf{Z}}_{A}$. Our aim is to study the fixed locus of $\iota_{A}$ following the degeneration argument used in [FMOS22] and [FMOS23]. One family of singular EPW cubes was studied in [Riz24]: their periods dominate the Heegner divisor $^{[3]}\mathscr{D}_{4,12}^{(2)}$ in the period domain. For the family that we construct below, the periods dominate another Heegner divisor, $^{[3]}\mathscr{D}_{4,2}^{(2)}$, in the same period domain (see [Riz24, Appendix A]) for the notation).
Let $(S,L)$ be a very general polarized K3 surface of degree $2$. The polarization $L$ induces a double cover $\varphi\coloneqq\varphi_{L}\colon S\to{\bf P}^{2}$ branched over a smooth sextic $\Gamma\subseteq{\bf P}^{2}$, and defines an involution $j$ of $S$. We have
$\operatorname{NS}(S^{[3]})={\bf Z}L\oplus{\bf Z}\delta.$
By [BC22, Proposition 4.2 and Theorem 1.1], the closed movable and nef cones of $S^{[3]}$ coincide and, by [BM14a, Proposition 13.1], they are spanned by the classes $L$ and $2L-\delta$. The class $2L-\delta$, of square $4$ and divisibility $2$, is therefore nef and big, but nonample on $S^{[3]}$. Therefore, the extremal ray ${\bf R}_{>0}(2L-\delta)$ induces a projective contraction
(16) $\pi\colon S^{[3]}\longrightarrow X.$
We show in Lemma 4.1 that this contraction is divisorial and describe its exceptional divisor $\Delta\subseteq S^{[3]}$.
Let $\bar{L}$ be the ample line bundle on the (singular) variety $X$ such that $\pi^{*}\bar{L}=L$. The biregular antisymplectic involution $j^{[3]}$ of $S^{[3]}$ induced by the involution $j$ of $S$ acts trivially on $\operatorname{NS}(S^{[3]})={\bf Z}L\oplus{\bf Z}\delta$ (and by $\operatorname{Id}$ on its orthogonal complement in $H^{2}(S^{[3]},{\bf Z})$), so it descends to an involution $\iota$ on $X$.
The data $S^{[3]}$, $L$, $\Delta$, and $j^{[3]}$ satisfy requirements (a)–(d) of [FMOS22, Section 2]. Hence, we can apply [FMOS22, Proposition 2.1 and Corollary 2.2] (as in [FMOS23, Section 3.1]) to obtain a family
(17) $f_{\mathscr{X}}\colon\mathscr{X}\longrightarrow D$
over a smooth pointed curve $(D,0)$. There is a relatively ample line bundle $\mathscr{L}$ on $\mathscr{X}$ and a fiberwise involution $\iota_{\mathscr{X}}$ of $\mathscr{X}$ such that:
  • the fiber over 0 is the variety $X$ with the ample line bundle $\bar{L}$ and involution $\iota$;
  • for any $t\in D\smallsetminus\{0\}$, $(\mathscr{X}_{t},\mathscr{L}_{t})$ is a smooth EPW cube $\widetilde{\mathsf{Z}}_{A_{t}}$ (for some Lagrangian $A_{t}\subseteq\textstyle{\bigwedge\thinspace^{3}}\thinspace{V}_{6}$) and $\iota_{t}$ is the natural EPW involution $\iota_{A_{t}}$.
In this section, we study the special fiber $(X,\iota)$ and the fixed locus of $\iota$.
4.1. The special fiber of $\mathscr{X}\to D$
The central fiber of the family $\mathscr{X}\to D$ is the singular variety $X$ obtained by the contraction (16) of $S^{[3]}$ defined by the nef and big extremal ray ${\bf R}_{\geq 0}(2L-\delta)$.
The variety $S^{[3]}$ is the moduli space $M_{S}(1,0,-2)$ of stable sheaves on $S$ with Mukai vector $v\coloneqq(1,0,-2)$. Recall that the vector $v$ lies in the Mukai lattice $\widetilde{\Lambda}=H^{0}(S,{\bf Z})\oplus H^{2}(S,{\bf Z})\oplus H^{0}(S,{\bf Z})$, and consider the orthogonal space $v^{\perp}\subseteq\widetilde{\Lambda}$. There is an identification
$\Theta\colon v^{\perp}\operatorname{\stackrel{{\scriptstyle{}_{\scriptstyle\sim}}}{{\longrightarrow}}}H^{2}(S^{[3]},{\bf Z})$
such that $L=\Theta(0,-L,0)$ and $\delta=\Theta(-1,0,-2)$ ([BM14a, Section 13]).
In [Bri08], Bridgeland showed that $S^{[3]}$ is also isomorphic to the moduli space $M_{\sigma}(v)$ of $\sigma$-stable objects in $D^{b}(S)$, where $\sigma$ is a stability condition contained in the distinguished connected component $\operatorname{Stab}^{\dagger}(S)$ (see [Bay18, Section 3] for the definition).
Contractions of moduli spaces of sheaves on K3 surfaces are described in [BM14a]. Their main result is that the wall and chamber decomposition of the movable cone of $M_{\sigma}(v)$ corresponds to the wall and chamber decomposition induced by $v$ on $\operatorname{Stab}^{\dagger}(S)$.
Moreover, the contraction associated to a certain wall $\mathscr{W}\subseteq\operatorname{Stab}^{\dagger}(S)$ is characterized via [BM14a, Theorem 5.7] in terms of the associated hyperbolic lattice $H_{\mathscr{W}}\in H^{*}_{\operatorname{alg}}(S,Z)$ (see [BM14a, Proposition 5.1] for the definition of $H_{\mathscr{W}}$). By definition, the hyperbolic lattice $H_{\mathscr{W}}$ contains $v$, and $\mathscr{W}$ corresponds to the wall $\Theta(H_{\mathscr{W}}^{\perp}\cap v^{\perp})$ of the movable cone of $M_{\sigma}(v)$.
In particular, in our case, contractions of $S^{[3]}$ are given by the two extremal rays ${\bf R}_{\geq 0}L$ and ${\bf R}_{\geq 0}(2L-\delta)$ of $\operatorname{Nef}(S)$. The extremal ray ${\bf R}_{\geq 0}L$ is orthogonal to the isotropic class $w\coloneqq(0,0,-1)$, which satisfies $w\cdot v=1$. The induced contraction is the Hilbert–Chow morphism $S^{[3]}\to S^{(3)}$ ([BM14a, Theorem 5.7(a)]).
The next lemma describes the contraction $\pi$ induced by the other extremal ray ${\bf R}_{\geq 0}(2L-\delta)$.
Lemma 4.1.
Let $(S,L)$ be a polarized K3 surface of degree $2$. The extremal ray ${\bf R}_{\geq 0}(2L-\delta)$ induces a divisorial contraction $\pi\colon S^{[3]}\to X$ with smooth exceptional divisor
$\Delta=\{Z\in S^{[3]}\mid Z\textnormal{ is contained in some curve }C\in|L|\}.$
Moreover, the restriction of $\pi$ to $\Delta$ is a ${\bf P}^{1}$-fibration
$\pi|_{\Delta}\colon\Delta\longrightarrow T,$
where the smooth hyper-Kähler fourfold $T$ is the moduli space of stable sheaves on $S$ with Mukai vector (0,L,4)(0,L,-4).
Proof.
We use the notation of [BM14a]. Under the isomorphism $\Theta$, the class $2L-\delta$ is the image of the vector w=(1,2L,2)w=(1,-2L,2). In particular, the hyperbolic lattice associated to the contraction induced by $2L-\delta$ is $H=\langle v,(1,-L,2)\rangle=w^{\perp}\subseteq H^{2}_{\operatorname{alg}}(S,{\bf Z})$. It is a lattice with intersection form $\begin{pmatrix}4&0\\ 0&-2\end{pmatrix}$, and contains the spherical class s=(1,L,2)s=(1,-L,2).
The corresponding potential wall $\mathscr{W}=\mathscr{W}_{H}$ inside the connected component $\operatorname{Stab}^{\dagger}(S)\subseteq\{(\alpha,\beta)\in{\bf R}_{>0}\times{\bf R}\}$ is a connected component of the (semi)-circular wall cut out by the equation
(18) $(\beta+2)^{2}+\alpha^{2}=2.$
Note that, by [Bri08, Lemma 6.2], every $(\alpha,\beta)\in{\bf R}_{>0}\times{\bf R}\smallsetminus\{(1,-1\}$ that satisfies Equation (18) defines a stability condition, hence the wall $\mathscr{W}$ is equal to
(19) $\mathscr{W}=\{\sigma_{\alpha,\beta}\mid(\beta+2)^{2}+\alpha^{2}=2,\ \alpha\in{\bf R}_{>0}\times(-\infty,-1)\}.$
We analyze the wall $\mathscr{W}$ using [BM14a, Theorem 5.7]. We first remark that $H$ contains no isotropic classes: indeed a vector $u=xv+ys\in H$, with $x,y\in{\bf Z}$, has square $4x^{2}-2y^{2}$ which is never 0.
Moreover, the wall $\mathscr{W}$ is not totally semistable. Since $\mathscr{W}$ is nonisotropic, to check this, we only need to show that there are no effective spherical classes $u$ with $(u,v)<0$. A spherical class $u\in H$ is a vector $u=xv+ys$ with $x,y\in{\bf Z}$ such that $2x^{2}-y^{2}=-1$, and the condition $(u,v)<0$ is equal to $x<0$. By [BM14a, Proposition 5.5], a class $u$ is effective if and only if $\mathop{\rm Re}\nolimits\left(\frac{Z_{\alpha,\beta}(u)}{Z_{\alpha,\beta}(v)}\right)>0$ for one (hence for all) $(\alpha,\beta)\in\mathscr{W}$. In particular, we can choose $\beta=-2$ and $\alpha=\sqrt{2}$.
Using [Bay18, Equation (3)], we compute $Z_{\alpha,\beta}(v)=-2\beta(\beta+2+i\alpha)$ and $Z_{\alpha,\beta}(s)=-2(\beta+1)(\beta+2+i\alpha)$ for $(\alpha,\beta)\in\mathscr{W}$. We obtain
$\frac{Z_{\sqrt{2},-2}(u)}{Z_{\sqrt{2},-2}(v)}=x+y\frac{Z_{\sqrt{2},-2}(s)}{Z_{\sqrt{2},-2}(v)}=x+\frac{y}{2}.$
Note that for each pair $(x,y)$ with $x<0$ that satisfies $2x^{2}-y^{2}=-1$, we have $x+\frac{y}{2}<0$, thus there are no effective spherical classes $u$ with $(u,v)<0$.
Therefore the wall $\mathscr{W}$ is nonisotropic not totally semistable, and the lattice $H$ contains the spherical class $s$, which is effective on $\mathscr{W}$ since $\frac{Z_{\sqrt{2},-2}(s)}{Z_{\sqrt{2},-2}(v)}=1/2$. By [BM14a, Theorem 5.7(a)], the associated contraction $\pi$ is divisorial of Brill–Noether type. In particular, by [BM14a, Lemma 7.1], $\pi$ contracts a divisor $\Delta$ of class $\Theta(s)=L-\delta$. Moreover, since $s$ is the Mukai vector of the line bundle $\mathscr{O}_{S}(-L)$, the divisor $\Delta$ is given by the condition $\operatorname{Hom}(\mathscr{O}_{S}(-L),\bullet)\neq 0$. In particular, this corresponds to length-3 subschemes $Z\subseteq S$ that are contained in some curve $C\in|L|$.
More precisely, following the proof of [BM14a, Lemma 7.4], the image $T\coloneqq\pi(\Delta)$ is isomorphic to the moduli space $M_{\sigma_{0}}(a)$, where $a\coloneqq v-s=(0,L,-4)$ and $\sigma_{0}=\sigma_{\alpha_{0},\beta_{0}}$ is a stability condition lying on the wall $\mathscr{W}$. By [BM14b, Example 9.7], we have equality $M_{\sigma_{0}}(a)=M_{S}(a)$ (by choosing $A=1$, $B=\beta_{0}$), hence $T$ is a smooth hyper-Kähler fourfold that parametrizes sheaves of the form $i_{C,*}(\xi)$, where $i_{C}\colon C\hookrightarrow S$ is the inclusion of a curve $C\in|L|$ in $S$ and $\xi$ is a torsion-free rank $1$ sheaf on $C$ of degree 3-3.
Finally, the fiber of $\pi$ over a point $A]\in M_{S}(a)$ is isomorphic to ${\bf P}(\operatorname{Ext}^{1}(\mathscr{O}_{S}(-L),A))$. Since
$\mathrm{ext}^{1}(\mathscr{O}_{S}(-L),A)=\langle s,v-s\rangle=-s^{2}=2,$
the map $\pi|_{E}\colon E\to M_{S}(a)$ is a ${\bf P}^{1}$-fibration.
If $C\in|L|$ is smooth (of genus 2), a torsion-free rank $1$ sheaf on $C$ of degree 3-3 is of the form $\mathscr{O}_{C}(-Z)$, with $Z\in C^{(3)}$. If $A\coloneqq i_{C,*}(\mathscr{O}_{C}(-Z))$, the preimage $\pi^{-1}([A])$ is equal to $|\mathscr{O}_{C}(Z)|$ (they have the same dimension by Riemann–Roch, and any degree $3$ effective divisor on $C$ equivalent to $Z$ defines a point in $S^{[3]}$ that is sent to $[A]$ via $\pi$). ∎
We keep the notation of the proof. Let $\sigma_{0}$ be a stability condition on the wall $\mathscr{W}$ defined in Equation (19). The moduli space $M_{\sigma_{0}}(v)$ has been constructed as a good moduli space in the sense of Alper (see [Taj23, Theorem 3.3]). By [BM14b, Proposition 8.1], there exists a morphism
(20) $f\colon X\longrightarrow M_{\sigma_{0}}(v)$
which is the identity on $X\smallsetminus\pi(\Delta)\simeq M_{\sigma}(v)\smallsetminus\Delta$ to $[F]$, and that sends $F]\in M_{S}(a)\simeq\pi(\Delta)$ to the polystable object $\mathscr{O}_{S}(H)\oplus F]$.
In particular, the morphism $f$ is a bijection of sets. We will show in Section 4.3 that $M_{\sigma_{0}}(v)$ is normal; this will imply that $f$ is an isomorphism by Zariski’s Main Theorem.
4.2. Fixed locus on the special fiber
We determine the set-theoretic fixed locus of the involution $\iota$ of $X$. We still denote by $\Gamma\subseteq S$ the ramification locus of $\varphi$, that is, the fixed curve of $j$.
Proposition 4.2.
The fixed locus of the involution $\iota$ of $X$ has two connected components:
$\operatorname{Fix}(\iota)={F_{2}}\sqcup{F_{3}},$
where
  • $F_{2}$} is irreducible of dimension 2 and contained in the fourfold $T=\pi(\Delta)$;
  • $F_{3}$} is irreducible of dimension 3 and its normalization is given by the restriction
    $\pi|_{\Gamma^{(3)}}\colon\Gamma^{(3)}\longrightarrow{F_{3}},$
    which is an étale double cover on the smooth divisor $\Delta\cap\Gamma^{(3)}\subseteq\Gamma^{(3)}$.
Proof.
We start by computing the fixed locus of $j^{[3]}$ on $S^{[3]}$. Since $S^{[3]}$ is a smooth hyper-Kähler variety and $j^{[3]}$ is an antisymplectic regular involution, $\operatorname{Fix}(j^{[3]})$ is smooth Lagrangian. In particular, each component has dimension $3$.
Clearly, $\Gamma^{[3]}\subseteq S^{[3]}$ is fixed by $j^{[3]}$, hence it is one connected component of $\operatorname{Fix}(j^{[3]})$. Therefore we can write
$\operatorname{Fix}(j^{[3]})=\Gamma^{[3]}\sqcup F^{\prime}.$
Note that $F^{\prime}$ is contained in the exceptional divisor $\Delta$. Indeed, if $Z\in S^{[3]}$ is fixed by $j^{[3]}$, then its support $\operatorname{Supp}(Z)$ is either contained in $\Gamma$, or of the form $\operatorname{Supp}(Z)=\{x+j(x)+y\}$. In particular, if $Z\notin\Gamma^{[3]}$ is fixed by $j^{[3]}$, then $\varphi(Z)$ is supported on two points, hence lies on a line in ${\bf P}^{2}$, so that $Z\in\Delta$.
This characterizes the fixed locus of $\iota$ outside the smooth fourfold $W=\pi(\Delta)\subseteq X$. We study the action of $\iota$ on $T=M_{S}(0,L,-4)$.
Lemma 4.3.
The action of the involution $\iota$ on the smooth fourfold $T=M_{S}(0,L,-4)$ is given by
(21) $\iota(i_{C,*}(\xi_{3}))=i_{C,*}(\omega_{C}^{\otimes-3}\otimes\xi_{3}^{\vee}),$
for any $C\in|L|$ and any degree 3-3 torsion-free rank 1 sheaf $\xi_{3}$ on $C$.
Proof.
Suppose that $C$ is smooth and $\xi_{3}=\mathscr{O}_{C}(-Z)$, where $Z=p+q+r\in C^{(3)}$. Then it satisfies $\iota\left(i_{C,*}\left(\mathscr{O}_{C}(-Z)\right)\right)=i_{C,*}\left(\mathscr{O}_{C}\left(-j^{[3]}(Z)\right)\right)$. By the adjunction formula, the canonical bundle of $C$ is equal to $\mathscr{O}_{C}(L)$. Since $\varphi|_{C}\colon C\to{\bf P}^{1}$ is a double cover, for any point $p\in C$, we have isomorphisms
$\omega_{C}\simeq\mathscr{O}_{C}(L)\simeq\varphi|_{C}^{*}(\mathscr{O}_{{\bf P}^{1}}(\varphi(p)))\simeq\mathscr{O}_{C}(p+j(p)).$
Therefore, we can rewrite
$\mathscr{O}_{C}\left(-j^{[3]}(Z)\right)=\left(\mathscr{O}_{C}(-j(p)-j(q)-j(r))\right)=\omega_{C}^{\otimes-3}\otimes\mathscr{O}_{C}(p+q+r).$
In particular, formula (21) is verified for any element of $T$ of the form $i_{C,*}(\mathscr{O}_{C}(-p-q-r))$ where $p,q,r$ are points on $C$ and $C\in|H|$ is a smooth curve.
By Bertini’s theorem, these elements form a dense open subset of $T$, therefore the involution on $T$ is given by formula (21) everywhere. ∎
Note that upon tensoring by $\mathscr{O}_{S}(L)$, we obtain an isomorphism
$T\simeq M_{S}(0,L,0),$
where elements of $M_{S}(0,L,0)$ are isomorphism classes of sheaves $i_{C,*}(\xi)$, where $i_{C}\colon C\to S$ is the inclusion of a curve $C\in|L|$ and $\xi$ is a torsion-free sheaf of rank $1$ and degree $1$ on $C$.
Under this isomorphism, the involution $\iota$ on $T$ corresponds to the involution $\iota_{0}$ on $M_{S}(0,L,0)$ given by
$\iota_{0}(i_{C,*}(\xi))=i_{C,*}(\omega_{C}\otimes\xi^{\vee}),$
which is one of the involutions studied in [FMOS22].
Using the description of the fixed locus of $\iota_{0}$ in [FMOS22, Proposition 4.2], we obtain that the fixed locus of $\iota|_{T}$ has two connected components $W_{+}$ and $W_{-}$, both smooth irreducible and 2-dimensional, which are the closures of the loci of $i_{C,*}(\xi\otimes\mathscr{O}(-L))$ where $\xi$ is an even, or odd, theta characteristic on a smooth curve $C$. Since $C$ has genus $2$, a theta characteristic $\xi$ on $C$ is odd if and only if $h^{0}(C,\xi)=1$ (the 6 odd theta characteristics correspond to the 6 Weierstrass points $x_{1},\dots,x_{6}$ of $C$; the 10 even theta characteristics correspond to the $x_{i}+x_{j}+x_{k}-g^{1}_{2}$, where $i,j,k\in\{1,\dots,6\}$ are distinct—this description counts each of them twice, using $x_{1}+\cdots+x_{6}=3g^{1}_{2}$).
To conclude, we show that $\operatorname{Fix}(\iota)$ is equal to the image $\pi(\operatorname{Fix}(j^{[3]}))$. Clearly, we have an inclusion $\pi(\operatorname{Fix}(j^{[3]}))\subseteq\operatorname{Fix}(\iota)$ and they are equal on $X\smallsetminus T$. We restrict ourselves to $T$, where $\operatorname{Fix}(\iota)\cap T=\operatorname{Fix}(\iota|_{T})=W_{+}\sqcup W_{-}$ and
$\pi\left(\operatorname{Fix}(j^{[3]})\right)\cap W=\pi\left(\operatorname{Fix}(j^{[3]})\cap\Delta\right)=\pi\left(\Gamma^{(3)}\cap\Delta\right)\cup\pi(F^{\prime}).$
Now, $\pi(F^{\prime})$ is contained in $W_{+}\sqcup W_{-}$, which has dimension 2. This implies that $F^{\prime}$ is contracted by $\pi$. We show that $\pi(F^{\prime})=W_{-}$. By definition of $W_{-}$, it is enough to show that for $i_{C,*}(\xi_{3})\in\pi(F^{\prime})$ general, the element $\xi=\xi_{3}\otimes\mathscr{O}_{C}(L)$ is an odd theta characteristic. The general element of $\pi(F^{\prime})$ is $\iota_{C,*}\mathscr{O}_{C}(-x-j(x)-y)\in T$, where $C$ is the pullback of the line through $\varphi(x)$ and $\varphi(y)$. Therefore, we obtain $\xi=\mathscr{O}_{C}(-x-j(x)-y)\otimes\mathscr{O}_{C}(x+j(x))=\mathscr{O}_{C}(-y)$, which is an odd theta characteristic by the discussion above, since $y$ belongs to $C\cap\Gamma$. Therefore, the morphism $\pi$ restricts to a ${\bf P}^{1}$-fibration $F^{\prime}\to W_{-}$.
The component $\Gamma^{(3)}$ intersects $\Delta$ in a divisor and is disjoint from $F^{\prime}$, hence $\pi(\Gamma^{(3)}\cap\Delta)$ is equal to $W_{+}$. The restriction of $\pi$ to $\Gamma^{3}\cap\Delta$ is finite, since the fiber over $i_{C,*}(\xi)$ is contained in $C^{(3)}\cap\Gamma^{(3)}\subseteq(C\cap\Gamma)^{3}$ Moreover, for any element $F=i_{C,*}(-Z)\in W_{+}$, for a smooth curve $C$, the intersection $\pi^{-1}(F)\cap\Gamma^{(3)}=|\mathscr{O}_{C}(-Z)|\cap\Gamma^{(3)}$ is given by $2$ points, namely $Z$ and the residual intersection $Z^{\prime}\coloneqq\Gamma\cdot C-Z$. Therefore, the induced morphism
(22) $\Gamma^{(3)}\cap\Delta\xrightarrow{\ 2:1\ }W_{+}$
is a double cover, and the corresponding involution sends collinear points $x+y+z\in\Gamma^{(3)}$ on a line $\ell$ to the residual intersection $\Gamma\cap\ell-(x+y+z)$. A general sextic $\Gamma\subseteq{\bf P}^{2}$ has no tritangent lines, so this involution is base-point-free. The double cover (22) is étale and, since the surface $W_{+}$ is smooth, so is $\Gamma^{(3)}\cap\Delta$.
Therefore, $\pi\left(\operatorname{Fix}(j^{[3]})\right)\cap T=\operatorname{Fix}(\iota|_{T})$. In particular, the fixed locus of $\iota$ is the image of the fixed locus of $j^{[3]}$.
This ends the proof of the proposition, with $F_{2}\coloneqq W_{-}$ and $F_{3}\coloneqq\pi(\Gamma^{[3]})$. ∎
Remark 4.4.
The scheme $F_{3}$ is semi-log-canonical (see [Kol23, Chapter 3, Definition 12]). Indeed, its normalization is the smooth scheme $\Gamma^{(3)}$ and the normalization morphism
$\pi\colon\Gamma^{(3)}\longrightarrow F_{3},$
restricts to an étale double cover over the divisor $E\subseteq\Gamma^{(3)}$ of collinear points of $\Gamma^{(3)}$.
4.3. Local description of $X$
Following [FMOS23], we give a local description of the variety $X$ and the involution $\iota$. Moreover, we show that the schematic fixed locus $\operatorname{Fix}(\iota)$ is a reduced local complete intersection.
The first step is to show that the morphism $f\colon X\to M_{\sigma_{0}}(v)$ constructed in Equation (20) is an isomorphism. We have already noted that $f$ is a bijection hence, by Zariski’s Main Theorem, it is enough to show that $M_{\sigma_{0}}(v)$ is normal.
By the analysis in Section 4.1, singular points of $M_{\sigma_{0}}(v)$ are given by $S$-equivalence classes of polystable sheaves of the form $E=\mathscr{O}(-L)\oplus F$, where $F$ is a stable vector sheaf in the fourfold $M_{S}(0,L,-4)$.
We study the local structure of $M_{\sigma_{0}}(v)$ at $[E]$ using the Kuranishi map. Consider the map
$\mu_{2}\colon\operatorname{Ext}^{1}(E,E)\longrightarrow\operatorname{Ext}^{2}(E,E)$
given by the Yoneda product. By [AS25, Corollary 4.1], there is an isomorphism of germs
(23) $(M_{\sigma_{0}}(v),[E])\simeq(\mu_{2}^{-1}(X)/\!\!/G,[0]),$
where the group $G$ is equal to $\operatorname{Aut}(\mathscr{O}(-L))\times\operatorname{Aut}(F)$ and acts on $\operatorname{Ext}^{1}(E,E)$ by conjugation.
The situation is analogous to that of [FMOS23, Remark 2.7 and Lemma 5.4]. Using their methods, we show the following lemma.
Lemma 4.5.
Let $[E]$ be a singular point of $M_{\sigma_{0}}(v)$. There is an isomorphism of analytic germs
$(X,[E])\simeq({\bf A}^{4}_{\bf C}\times Q,0)$
where $Q=\operatorname{Spec}\left({\bf C}[u_{1},u_{2},u_{3}]/u_{1}^{2}-u_{2}u_{3}\right)$.
In particular, the good moduli space $M_{\sigma_{0}}(v)$ is normal.
Proof.
Up to changing the singular point $[E]$ in its $S$-equivalence class, we can suppose that $E$ is of the form $T\oplus F$, where $T$ is the stable spherical sheaf $\mathscr{O}(-L)$ and $F$ is stable with Mukai vector a=(0,L,4)a=(0,L,-4). By Equation (23), the germ of $X$ at the singular point $E$ is equal to $\mu_{2}^{-1}(0)/\!\!/G$.
We describe the morphism $\mu_{2}$ on $\operatorname{Ext}^{1}(E,E)$. Since $T$ is spherical, we have decompositions
$\operatorname{Ext}^{1}(E,E)=\operatorname{Ext}^{1}(F,F)\oplus\operatorname{Ext}^{1}(T,F)\oplus\operatorname{Ext}^{1}(F,T),$
and
$\operatorname{Ext}^{2}(E,E)=\operatorname{Ext}^{2}(F,F)\oplus\operatorname{Ext}^{2}(T,T)\simeq{\bf C}^{2},$
where we used that $\operatorname{Ext}^{2}(T,F)=\operatorname{Ext}^{2}(F,T)=0$, since $T$ and $F$ are nonisomorphic stable sheaves of same phase. The Yoneda product $\mu_{2}$ is given by
$(\eta,a,b)\longmapsto(\eta^{2}+b\circ a,a\circ b).$
Note that, since $F$ is a stable point of the moduli space $M_{\sigma_{0}}(a)$, the Kuranishi map for $F$ is trivial, hence $\eta^{2}=0$ for any $\eta\in\operatorname{Ext}^{1}(F,F)$. Additionally, the trace map defines isomorphisms $\operatorname{Ext}^{2}(F,F)\simeq{\bf C}$ and $\operatorname{Ext}^{2}(T,T)\simeq{\bf C}$ such that $a\circ b=-b\circ a$ for any $(a,b)$ in $\operatorname{Ext}^{1}(T,F)\oplus\operatorname{Ext}^{1}(F,T)$.
Therefore, the kernel of $\mu_{2}$ is equal to
$\mu_{2}^{-1}(0)=\operatorname{Ext}^{1}(F,F)\times\{(a,b)\in\operatorname{Ext}^{1}(T,F)\oplus\operatorname{Ext}^{1}(F,T)\mid a\circ b=0\}.$
The vector space $\operatorname{Ext}^{1}(F,F)$ has dimension $a^{2}+2=4$, and the spaces $\operatorname{Ext}^{1}(T,F)$ and $\operatorname{Ext}^{1}(F,T)$ have dimension $a\cdot(v-a)=2$. Given a basis $(a_{1},a_{2})$ of $\operatorname{Ext}^{1}(T,F)$, let $(b_{1},b_{2})$ be a basis of $\operatorname{Ext}^{2}(F,T)$ such that $a_{1}\circ b_{2}=a_{2}\circ b_{1}=0$. We obtain an isomorphism
$\mu_{2}^{-1}(0)\simeq{\bf A}_{\bf C}^{4}\times\operatorname{Spec}\left({\bf C}[a_{1},a_{2},b_{1},b_{2}]/(a_{1}b_{1}+a_{2}b_{2})\right).$
Now the action of the group $G=\operatorname{Aut}(T)\oplus\operatorname{Aut}(F)\simeq({\bf C}^{*})^{2}$ is given by conjugation on elements of $\operatorname{Ext}^{1}(E,E)$. Therefore, it is trivial on $\operatorname{Ext}^{1}(F,F)$ and $(g,h)\in({\bf C}^{*})^{2}$ acts as the multiplication by $gh^{-1}$ on $\operatorname{Ext}^{1}(T,F)$ and as the multiplication by $hg^{-1}$ on $\operatorname{Ext}^{1}(F,T)$. Therefore, as in [FMOS23, Remark 2.7], the quotient $\mu_{2}^{-1}(0)/\!\!/G$ is equal to ${\bf A}_{\bf C}^{4}\times Q$, where
$Q=\operatorname{Spec}\left({\bf C}[u_{1},u_{2},u_{3}]/(u_{1}^{2}-u_{2}u_{3})\right),$
where $u_{1}=a_{1}b_{1}$, $u_{2}=-a_{1}b_{2}$ and $u_{3}=a_{2}b_{1}$.
This shows in particular that $M_{\sigma_{0}}(v)$ is a local complete intersection. Since its singular locus is parametrized by $M_{S}(v)$, it has codimension 2. Therefore the moduli space $M_{\sigma_{0}}(v)$ is normal. ∎
As already noted, the above lemma implies the wanted modular description for the special fiber of $\mathscr{X}\to B$.
Corollary 4.6.
The morphism $f$ given in Equation (20) is an isomorphism, hence the central fiber $X$ of the family $\mathscr{X}\to B$ is isomorphic to $M_{\sigma_{0}}(v)$.
This allows us to determine the local structure of the schematic fixed locus $\operatorname{Fix}(\iota)$.
Lemma 4.7.
Let $[E]$ be a singular point of $X$ and consider the isomorphism of germs
$(X,[E])\simeq({\bf A}^{4}_{\bf C}\times Q,0)$
from Lemma 4.5. If the point $[E]$ lies on the component $F_{3}$ or $\operatorname{Fix}(\iota)$ (see Proposition 4.2), this isomorphism restricts to an isomorphism of germs
$(\operatorname{Fix}(\iota),[E])\simeq({\bf A}^{2}_{\bf C}\times B,0),$
where $B=\operatorname{Spec}\left({\bf C}[u_{2},u_{3}]/(u_{2}u_{3})\right)$. In particular, $\operatorname{Fix}(\iota)$ is reduced at points of $F_{3}$.
Proof.
By Corollary 4.6, the variety $X$ is isomorphic to $M_{\sigma_{0}}(v)$, and this gives the first isomorphism of germs. We go back to the notation of the proof of Lemma 4.5.
We start by describing the action of the involution on the germ ${\bf A}_{\bf C}^{4}\times Q$. Since $[E]$ is fixed by $\iota$, the factor $F\in M_{S}(a)$ is fixed by the restriction of $\iota$ to $M_{S}(a)=W\subseteq X$ (see Section 4.1). Since $M_{S}(a)$ is a smooth hyper-Kähler fourfold and $\iota$ restricts to an antisymplectic involution on it, its fixed locus is smooth of dimension 2. Hence, on germs, the fixed locus of $\iota|_{{\bf A}^{4}_{\bf C}}$ is equal to ${\bf A}_{\bf C}^{2}$.
We now analyze the action of $\iota$ on $Q$. Note that the fixed eigenspace for the action of $\iota$ on $\operatorname{Ext}^{1}(T,F)$ has dimension $1$. Indeed, since ${\bf P}(\operatorname{Ext}^{1}(T,F))$ coincide with the fiber of $\pi$ over the point $E$ (see Lemma 4.3), this is equivalent to say the fixed locus of $j^{(3)}|_{{\bf P}(\operatorname{Ext}^{1}(T,F))}$ is finite (when $F$ is $j^{(3)}$-invariant), which follows from the proof of Proposition 4.2.
Moreover the action of $\iota$ on $\operatorname{Ext}^{1}(F,T)$ is dual to the one on $\operatorname{Ext}^{1}(T,F)$, therefore the action of $\iota$ on $Q$ is the one described in [FMOS23, Proposition 5.4]. In particular, its fixed locus is equal to $B=\operatorname{Spec}({\bf C}[u_{2},u_{3}]/u_{2}u_{3})$.
This implies that the schematic fixed locus $\operatorname{Fix}(\iota)$ is reduced. Indeed, it is smooth at any smooth point of $X$, since the action $\iota$ restricts to a regular action on the smooth locus of $X$, and the reducedness at singular points follows from its local description, since $B$ is reduced. ∎
Remark 4.8.
A similar argument shows that the whole fixed locus of $\iota$ is reduced.
5. Degeneration of ${\mathsf{W}}_{A}$
Given the family $\mathscr{X}\to D$ studied in Section 4, we consider the schematic fixed locus $\operatorname{Fix}(\iota_{\mathscr{X}})\to\mathscr{D}$ of $\iota_{\mathscr{X}}$ on $\mathscr{X}$ (see the proof of [FMOS23, Proposition 5.6] for the definition of the schematic structure on the fixed locus). For $t\neq 0$, the fiber $\operatorname{Fix}(\iota_{\mathscr{X}})_{t}=\operatorname{Fix}(\iota_{t})$ is the smooth Lagrangian threefold ${\mathsf{W}}_{A_{t}}$. The reducedness of the fixed locus of $\iota$ studied in Section 4.3 implies $\operatorname{Fix}(\iota_{\mathscr{X}})_{0}=\operatorname{Fix}(\iota)=F_{3}\cup F_{2}$.
Let $\mathscr{X}^{*}$ be the pullback of $\mathscr{X}\to D$ to $D\smallsetminus\{0\}$. The family $\mathscr{X}^{*}$ is smooth, hence the fixed locus $\operatorname{Fix}_{\mathscr{X}^{*}}(\iota_{\mathscr{X}^{*}})$ of $\iota_{\mathscr{X}^{*}}$ is also smooth and its fibers are equal to $\operatorname{Fix}(\iota_{t})$ for $t\neq 0$. Let $\mathscr{Y}$ be the closure of $\operatorname{Fix}_{\mathscr{X}^{*}}(\iota_{\mathscr{X}^{*}})$ inside $\mathscr{X}$, with induced morphism
(24) $f\colon\mathscr{Y}\longrightarrow D.$
The morphism $f$ is flat by [Liu02, 4.3 Proposition 3.9] since it is dominant and $\mathscr{Y}$ is integral. Therefore, it is equidimensional and $\mathscr{Y}_{0}$ has dimension $3$. Moreover, $\mathscr{Y}$ is contained in $\operatorname{Fix}(\iota_{\mathscr{X}})$, hence $\mathscr{Y}_{0}$ is equal to $F_{3}$ (scheme-theoretically).
Theorem 5.1.
Let $A\in\operatorname{\mathsf{LG}}(\textstyle{\bigwedge\thinspace^{3}}\thinspace{V}_{6})\smallsetminus\Sigma\cup\Gamma$ be a Lagrangian, with associated smooth EPW cube $\widetilde{\mathsf{Z}}_{A}$. The external Hodge numbers of ${\mathsf{W}}_{A}$ can be computed as
$h^{p,0}({\mathsf{W}}_{A})=h^{p}(F_{3},\mathscr{O}_{F_{3}}).$
Proof.
Hodge numbers are invariant under smooth deformations, hence it is enough to compute $h^{p,0}({\mathsf{W}}_{A})$ when ${\mathsf{W}}_{A}$ is a smooth fiber of the morphism $f\colon\mathscr{Y}\to D$ defined in (24).
All fibers of $f$ are slc: the central fiber is isomorphic to $F_{3}$, which is slc by Remark 4.4, and the other fibers ${\mathsf{W}}_{A}$ are smooth. Therefore, since $f$ is flat and proper, [Kol23, Corollary 2.64] implies that $h^{p}(\mathscr{Y}_{t},\mathscr{O}_{\mathscr{Y}_{t}})$ is independent of $t\in D$. The theorem follows. ∎
5.1. Cohomology groups $H^{p}(F_{3},\mathscr{O}_{F_{3}})$
We study the cohomology of the structure sheaf of $F_{3}$. We keep the notation of Section 4.2.
Let $\Gamma\subseteq{\bf P}^{2}$ be a smooth plane sextic and let $E\subseteq\Gamma^{(3)}$ be the (smooth) divisor of collinear 3-points on $\Gamma$, with involution $\sigma$ defined by
$\sigma(x+y+z)=(C\cdot\ell)-x-y-z,$
for any $x+y+z\in E$ that lies on a line $\ell$.
By Proposition 4.2, the scheme $F_{3}$ is obtained from $\Gamma^{(3)}$ via the morphism
(25) $\eta\colon\Gamma^{(3)}\longrightarrow F_{3}$
given by the restriction of the contraction $\pi\colon S^{[3]}\to X$ to $\Gamma^{(3)}\subseteq S^{[3]}$. The morphism $\eta$ is an isomorphism on $\Gamma^{(3)}\smallsetminus E$, and $\eta|_{E}\colon E\to W\coloneqq\eta(E)$ is the (étale) double cover induced by the involution $\sigma$ defined above. Notice that $\eta$ is in particular the normalization of $F_{3}$.
Consider the morphism of sheaf $\mathscr{O}_{F_{3}}\to\eta_{*}\mathscr{O}_{\Gamma^{(3)}}$. It is injective, as $\eta$ is dominant. Moreover, since $\eta$ is an isomorphism outside $E$, the cokernel of $\eta$ is supported on $W$. Since the restriction $\eta|_{E}$ is a double cover, there is a decomposition
$(\eta|_{E})_{*}\mathscr{O}_{E}=\mathscr{O}_{W}\oplus\mathscr{R},$
where $\mathscr{O}_{W}$ and $\mathscr{R}$ are respectively the invariant and anti-invariant parts with respect to the action of $\sigma^{*}$. Since $\mathscr{O}_{F_{3}}|_{W}=\mathscr{O}_{W}$ and $\eta_{*}\mathscr{O}_{\Gamma^{(3)}}|_{W}=(\eta|_{E})_{*}\mathscr{O}_{E}$, we have a short exact sequence
(26) $0\longrightarrow\mathscr{O}_{F_{3}}\longrightarrow\eta_{*}\mathscr{O}_{\Gamma^{(3)}}\longrightarrow j_{*}\mathscr{R}\longrightarrow 0$
where $j$ is the embedding $W\hookrightarrow F_{3}$ and the map $\eta_{*}\mathscr{O}_{\Gamma^{(3)}}\to j_{*}\mathscr{R}$ is induced by taking the restriction to $W$ and projecting to $\mathscr{R}$. This induces a long exact sequence in cohomology
(27) $0$ $\to H^{1}(F_{3},\mathscr{O}_{F_{3}})\to H^{1}(\Gamma^{(3)},\mathscr{O}_{\Gamma^{(3)}})\xrightarrow{\ \beta_{1}\ }H^{1}(W,\mathscr{R})\to H^{2}(F_{3},\mathscr{O}_{F_{3}})\to$
$\to H^{2}(\Gamma^{(3)},\mathscr{O}_{\Gamma^{(3)}})\xrightarrow{\ \beta_{2}\ }H^{2}(W,\mathscr{R})\to H^{3}(F_{3},\mathscr{O}_{F_{3}})\to H^{3}(\Gamma^{(3)},\mathscr{O}_{\Gamma^{(3)}})\to 0,$
where we used that $H^{3}(W,\mathscr{R})=0$ since $W$ has dimension $2$, and $H^{0}(W,\mathscr{R})=0$ because of the equality ${\bf C}=H^{0}(E,\mathscr{O}_{E})=H^{0}(W,(\eta|_{E})_{*}\mathscr{O}_{E})=H^{0}(W,\mathscr{O}_{W})$.
The map $\beta_{i}\colon H^{i}(\Gamma^{(3)},\mathscr{O}_{\Gamma^{(3)}})\to H^{i}(W,\mathscr{R})$ factors as
$H^{i}(\Gamma^{(3)},\mathscr{O}_{\Gamma^{(3)}})\xrightarrow{\ \mathsf{res}_{i}\ }H^{i}(E,\mathscr{O}_{E})=H^{i}(W,(\eta|_{E})_{*}\mathscr{O}_{E})\longrightarrow H^{i}(W,\mathscr{R}),$
where the morphism $\mathsf{res}_{i}$ is induced by the restriction to $E\subseteq\Gamma^{(3)}$ and the second map is induced by the projection $(\eta|_{E})_{*}\mathscr{O}_{E}\to\mathscr{R}$.
Lemma 5.2.
The restrictions
$\mathsf{res}_{0,q}\colon H^{0}(\Gamma^{(3)},\Omega_{\Gamma^{(3)}}^{q})\longrightarrow H^{0}(E,\Omega_{E}^{q})$
are injective for $q\in\{0,1,2\}$.
Proof.
Since $\Gamma$ is a smooth plane sextic curve, it is not trigonal, hence the symmetric product $\Gamma^{(3)}$ embeds into $\operatorname{Jac}(\Gamma)$ and the restrictions
[…]
are isomorphisms for $q\in\{0,1,2,3\}$ ([Mac62, (11.1)]). Therefore, it is enough to consider the restriction morphisms induced by the inclusion $E\subseteq\Gamma^{(3)}\subseteq\operatorname{Jac}(\Gamma)$. Moreover, since
[…]
it is enough to show the injectivity of the restriction
$r\colon H^{0}(\operatorname{Jac}(\Gamma),\Omega^{2}_{\operatorname{Jac}(\Gamma)})\longrightarrow H^{0}(E,\Omega^{2}_{E}).$
We show that the surface $E\subseteq\operatorname{Jac}(\Gamma)$ is nondegenerate in the sense of [Ran80, Section II], which implies by [Ran80, Lemma II.1] that $r$ is injective.
Let $\theta$ be the canonical principal polarization on $\operatorname{Jac}(\Gamma)$. By [Mac62, (16.2)], the cohomology class of $E$ in $H^{2}(\Gamma^{(3)},{\bf Z})$ is equal to
$6\eta+\theta|_{\Gamma^{(3)}}\in H^{2}(\Gamma^{(3)},{\bf Z}),$
where $\eta\in H^{2}(\Gamma^{(3)},{\bf Z})$ is the class of $\Gamma^{(2)}\subseteq\Gamma^{(3)}$. [^4] The divisor $E\subseteq\Gamma^{(3)}$ is not ample. Indeed by [Kou93, Lemma (1)], we have $(\theta-6\eta)^{3}=\sum_{i=0}^{3}{3\choose i}(-6)^{3-i}\frac{10!}{(10-i)!}=-36$ By the projection formula, the class of $E$ in $\operatorname{Jac}(\Gamma)$ is therefore given by
$E]=-6[\Gamma^{(2)}]+[\Gamma^{(3)}]\cdot\theta.$
Since $\Gamma^{(i)}]\in H^{*}(\operatorname{Jac}(\Gamma),{\bf Z})$ is equal to $\frac{\theta^{10-i}}{(10-i)!}$, it follows that $E]=2\theta^{8}/8!$ in $H^{*}(\operatorname{Jac}(\Gamma),{\bf Z})$. Therefore, $[E]$ is a positive multiple of $\theta^{8}$, hence $E$ is nondegenerate by [Ran80, Corollary II.2]. ∎
Proposition 5.3.
The cohomology groups of the structure sheaf of $F_{3}$ are
$H^{1}(F_{3},\mathscr{O}_{F_{3}})$ $=0,$
$H^{2}(F_{3},\mathscr{O}_{F_{3}})$ $=H^{2,0}(\Gamma^{(3)})\oplus\operatorname{Coker}(\mathsf{res}_{1}),$
$H^{3}(F_{3},\mathscr{O}_{F_{3}})$ $=H^{3,0}(\Gamma^{(3)})\oplus H^{2}(W,\mathscr{R}).$
Proof.
We start by showing that in the long exact sequence (27), the morphism $\beta_{1}$ is injective and $\beta_{2}$ is the zero map. We have a factorization
$\beta_{i}\colon H^{i}(\Gamma^{(3)},\mathscr{O}_{\Gamma^{(3)}})\hookrightarrow H^{i}(E,\mathscr{O}_{E})\twoheadrightarrow H^{i}(W,\mathscr{R}),$
where the first map is injective by Lemma 5.2 and Hodge theory, since $\Gamma^{(3)}$ and $E$ are smooth by Proposition 4.2, and the second map is the projection to the anti-invariant part of $H^{i}(E,\mathscr{O}_{E})$ with respect to the action of $\sigma^{*}$.
Consider the embedding $E\subseteq\Gamma^{(3)}\subseteq\operatorname{Jac}^{3}(\Gamma)$, as in the proof of Lemma 5.2. By (28) and (29), the involution $\sigma^{*}$ acts by multiplication by $(-1)^{q}$ on the subspace $H^{0}(\Gamma^{(3)},\Omega^{q}_{\Gamma^{(3)}})\subseteq H^{0}(E,\Omega^{q}_{E})$, hence also on $H^{q}(\Gamma^{(3)},\mathscr{O}_{\Gamma^{(3)}})\subseteq H^{q}(E,\mathscr{O}_{E})$.
In particular, $H^{1}(\Gamma^{(3)},\mathscr{O}_{\Gamma^{(3)}})$ is anti-invariant for the action of $\sigma^{*}$ on $H^{1}(E,\mathscr{O}_{E})$, hence is contained in $H^{1}(W,\mathscr{R})$. Therefore, the morphism $\beta_{1}$ is injective. Analogously, $H^{2}(\Gamma^{(3)},\mathscr{O}_{\Gamma^{(3)}})$ is all invariant for the action of $\sigma^{*}$, therefore the morphism $\beta_{2}$ is the zero map.
Therefore, the long exact sequence (27) gives $H^{1}(F_{3},\mathscr{O}_{F_{3}})=0$ and exact sequences
$0\to H^{1}(\Gamma^{(3)},\mathscr{O}_{\Gamma^{(3)}})\xrightarrow{\ \beta_{1}\ }H^{1}(W,\mathscr{R})\to H^{2}(F_{3},\mathscr{O}_{F_{3}})\to H^{2}(\Gamma^{(3)},\mathscr{O}_{\Gamma^{(3)}})\to 0$
and
$0\to H^{2}(W,\mathscr{R})\to H^{3}(F_{3},\mathscr{O}_{F_{3}})\to H^{3}(\Gamma^{(3)},\mathscr{O}_{\Gamma^{(3)}})\to 0,$
from which we deduce the wanted statements. ∎
Using Theorem 5.1, Proposition 3.3, (28), and (29), we obtain information on some Hodge numbers of the smooth threefold ${\mathsf{W}}_{A}$.
Corollary 5.4.
Assume that the EPW cube $\widetilde{\mathsf{Z}}_{A}$ is smooth. We have
$h^{0,1}({\mathsf{W}}_{A})$ $=0,$
$h^{0,2}({\mathsf{W}}_{A})$ $\geq 45,$
$h^{0,3}({\mathsf{W}}_{A})$ $=131+h^{0,2}({\mathsf{W}}_{A}),$
$h^{1,2}({\mathsf{W}}_{A})$ $=470+h^{0,2}({\mathsf{W}}_{A})+h^{1,1}({\mathsf{W}}_{A}).$
Corollary 5.5.
Assume that the EPW cube $\widetilde{\mathsf{Z}}_{A}$ is smooth. Its submanifold ${\mathsf{W}}_{A}$ is rigid in $\widetilde{\mathsf{Z}}_{A}$.
Proof.
Since ${\mathsf{W}}_{A}$ is Lagrangian, the normal sheaf $N_{{\mathsf{W}}_{A}/\widetilde{\mathsf{Z}}_{A}}$ is isomorphic to $\Omega_{{\mathsf{W}}_{A}}^{1}$. Therefore,
$h^{0}({\mathsf{W}}_{A},N_{{\mathsf{W}}_{A}/\widetilde{\mathsf{Z}}_{A}})=h^{0}({\mathsf{W}}_{A},\Omega_{{\mathsf{W}}_{A}}^{1})=h^{0,1}({\mathsf{W}}_{A})=0$
and the submanifold ${\mathsf{W}}_{A}$ does not deform in $\widetilde{\mathsf{Z}}_{A}$. ∎